OpenStudy (anonymous):

Write an equation of the line passing through (-3,2) and is parallel to the line x-y = 7?

OpenStudy (anonymous):

Try subbing (-3,2) as (x,y) in the formula y=mx+c.

OpenStudy (anonymous):

First find slope x-y=7 -y=-x+7 y=x-7 Slope is 1 use the equation \[y-y _{1}=m(x-x_{1})\] y-2=1(x+3) y-2=x+3 y=x+5

OpenStudy (anonymous):

Or -x+y=-5